OpenStudy (anonymous):

the formula is A=P(1+r)^2...A is for dollars that P dollars will grow into 2 years of interest rates r.. what interest rate will cause $1000 to grow to $1071.23 in 2 years.. am looking for interest rate..

OpenStudy (anonymous):

\[1071.23=1000(1+r)^2\] \[1.07123=(1+r)^2\] \[1+r=\sqrt{1.07123}\] \[r=\sqrt{1.07123}-1\]